# Sparse Covariance Neural Networks

### v1

@@ Abstract
Covariance Neural Networks (VNNs) perform graph convolutions on the covariance matrix of tabular data and achieve success in a variety of applications. However, the empirical covariance matrix on which the VNNs operate may contain many spurious correlations, making VNNs’ performance inconsistent due to these noisy estimates and decreasing their computational efficiency. To tackle this issue, we put forth Sparse coVariance Neural Networks (S-VNNs), a framework that applies sparsification techniques on the sample covariance matrix before convolution. When the true covariance matrix is sparse, we propose hard and soft thresholding to improve covariance estimation and reduce computational cost. Instead, when the true covariance is dense, we propose stochastic sparsification where data correlations are dropped in probability according to principled strategies. We show that S-VNNs are more stable than nominal VNNs as well as sparse principal component analysis. By analyzing the impact of sparsification on their behavior, we provide novel connections between S-VNN stability and data distribution. We support our theoretical findings with experimental results on various application scenarios, ranging from brain data to human action recognition, and show an improved task performance, stability, and computational efficiency of S-VNNs compared with nominal VNNs.
1 Introduction
@@
Covariance-based data processing is key to machine learning pipelines due to its ability to whiten data distributions, identify principal directions, and estimate the interdependencies among features. Such advantages have been shown in several applications including brain connectivity estimation (Bessadok et al., 2022; Qiao et al., 2016), financial data (de Miranda Cardoso et al., 2020; Wang & Aste, 2022) and human action recognition through motion sensors measurements (Liao et al., 2022; Wang et al., 2023). One prominent approach is Principal Component Analysis (PCA), which maximizes the variance of data points by projecting them onto the eigenvectors of their covariance matrix (Jolliffe & Cadima, 2016). However, PCA-based approaches are unstable to errors in the covariance estimation, i.e., a poor estimate of the covariance matrix and its eigenvectors may lead to unpredictably bad results, especially in high-dimensional and low-data regimes or when the covariance matrix eigenvalues are close to each other (Jolliffe, 2002; Paul, 2007; Baik et al., 2005). To overcome this issue, coVariance Neural Networks (VNNs) have been proposed (Sihag et al., 2022). VNNs apply graph convolutions to the covariance matrix, an operation that, similarly to PCA, learns the importance of the principal directions of the data. Due to their graph convolution operation, they inherit the stability property of GNNs (Gama et al., 2020b) and extend it to errors in covariance estimation, which guarantees more reliable performance even in the presence of poor covariance estimates since their output difference when operating on the sample and true covariance is bounded (Sihag et al., 2022; 2024; 2023). Nevertheless, VNNs operate on the sample covariance matrix, which results in two limitations. First, VNNs are sensitive to settings in which covariance estimation is problematic, especially in high-dimensional and low-data regimes (Paul, 2007; Baik et al., 2005; Féral & Péché, 2007). Second, VNNs are computationally expensive and memory inefficient as the sample covariance matrix is typically dense no matter whether the true covariance matrix is dense or sparse due to sample estimation errors, which is especially problematic on high-dimensional datasets.
@@
To overcome these limitations, we propose sparsification-based covariance regularizers for VNNs to improve the covariance matrix estimation. We perform stability analysis for sparse VNNs (S-VNNs) to investigate the effects of covariance sparsification techniques on the performance of VNNs. When true covariance matrices are sparse, we propose hard and soft thresholding strategies and show that S-VNNs are more robust to both the finite-data estimation error and the covariance sparsification thresholding. For generic scenarios where true covariance matrices are not necessarily sparse, we propose a stochastic sparsification approach and prove the stability of S-VNNs in this setting. Our findings quantify a trade-off between stability and sparsification (i.e., efficiency). That is, dropping covariances according to their absolute values leads to higher stability but, in some cases, low sparsification, whereas dropping a fixed percentage of covariance values improves sparsification at the expense of stability. We corroborate our findings with experiments on synthetic and real datasets highlighting the benefits of sparsity on covariance-based neural processing both in terms of performance and computation. Our specific contributions are as follows.
  1. (C1)
@@
    Sparse covariance neural networks. We propose covariance sparsification techniques for VNNs to improve performance and reduce computation. When the true covariance is sparse, we perform hard and soft covariance thresholding. When the true covariance is dense, we propose a stochastic sparsification framework that allows controlling the sparsification level, and design specific strategies in this framework based on application requirements.
  2. (C2)
@@
    Stability to covariance sparsification. We characterize the effects of covariance sparsification techniques on the VNN performance through stability analysis. We show that S-VNNs can not only reduce the computation but also improve the stability under appropriate settings, compared to the nominal VNN and the sparse PCA.
  3. (C3)
@@
    Empirical validation. We validate our theoretical findings with experiments on real and synthetic datasets, demonstrating the importance of sparsification in low-data settings and the improved efficiency of S-VNNs on brain data and human action recognition use cases.
2 Problem Formulation
Consider $t$ samples $\{\mathbf{x}_{i}\}_{i=1}^{t}$ of a random vector $\mathbf{x}\in\mathbb{R}^{N}$ with mean $\bm{\mu}=\mathbb{E}[\mathbf{x}]$, covariance $\mathbf{C}=\mathbb{E}[(\mathbf{x}-\bm{\mu})(\mathbf{x}-\bm{\mu})^{\mathsf{T}}]$ and respective estimates $\bm{\hat{\mu}}=1/t\sum_{i=1}^{t}\mathbf{x}_{i}$, $\mathbf{\hat{C}}=1/t\sum_{i=1}^{t}(\mathbf{x}_{i}-\bm{\hat{\mu}})(\mathbf{x}_{i}-\bm{\hat{\mu}})^{\mathsf{T}}$. The estimated covariance $\mathbf{\hat{C}}$ captures the data interdependencies and can be used for principal component analysis (PCA) via its eigendecomposition $\mathbf{\hat{C}}=\mathbf{\hat{V}}\mathbf{\hat{\Lambda}}\mathbf{\hat{V}}^{\mathsf{T}}$ with eigenvectors $\mathbf{\hat{V}}=[\mathbf{\hat{v}}_{0},\dots,\mathbf{\hat{v}}_{N-1}]\in\mathbb{R}^{N\times N}$ and eigenvalues $\mathbf{\hat{\Lambda}}=\text{diag}(\hat{\lambda}_{0},\dots,\hat{\lambda}_{N-1})$, to maximize the variance of the data points by projecting them on its eigenvectors, i.e., $\mathbf{\tilde{x}}=\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{x}$. However, this finite-sample PCA is unstable as the estimated eigenvectors $\mathbf{\hat{V}}$ may differ significantly from the true directions $\mathbf{V}$ especially when the true eigenvalues in $\mathbf{\Lambda}$ are close to each other (Jolliffe, 2002).
To overcome this issue, Sihag et al. (2022) proposed graph convolutional covariance filters as an extension of PCA that is stable to finite-sample estimation uncertainties even with close eigenvalues. Each data point is seen as a node of a graph and the data sample $\mathbf{x}$ as a graph signal, where the covariance $\mathbf{C}$ represents the graph structure. Covariance filters build on the standard graph convolutional filters, which shift a signal $\mathbf{x}$ over a graph represented by the shift operator $\mathbf{S}$ (e.g., adjacency matrix or Laplacian) according to the equation $\sum_{k=0}^{K}h_{k}\mathbf{S}^{k}\mathbf{x}$, where $K$ is the order of the filter and $h_{k}$ are learnable parameters (Gama et al., 2020c; Ortega et al., 2018; Isufi et al., 2024). Several popular GNNs, such as GCN (Kipf & Welling, 2016), are special instances of the generic graph convolutional filter. Building on this, the covariance filter of order $K$ and the VNN layer, which assembles a filter bank of $F_{\text{in}}\times F_{\text{out}}$ covariance filters and a non-linear activation function $\sigma(\cdot)$, are
$\mathbf{u}=\sum_{k=0}^{K}h_{k}\mathbf{\hat{C}}^{k}\mathbf{x}=\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}~{}~{}\textnormal{and}~{}~{}\mathbf{u}^{l}_{f}=\sigma\left(\sum_{g=1}^{F_{\text{in}}}\mathbf{H}^{l}(\mathbf{\hat{C}})\mathbf{u}^{l-1}_{g}\right)~{}f=1,\ldots,F_{\text{out}},~{}l=1,\ldots,L.$ (1)
This operation is related to PCA as it processes the eigenvectors of the covariance matrix (see Appendix A for details). We denote the VNN architecture as $\mathsf{\Phi}(\mathbf{x},\mathbf{\hat{C}},\mathcal{H})$, where =$\mathcal{H}=\{h_{klgf}\}$ contains all parameters. The output of the last layer $\mathbf{u}^{L}=\Phi(\mathbf{x},\mathbf{\hat{C}},\mathcal{H})$ contains the final representations generated by VNN and can be directly used for a downstream task (e.g., classification or regression) or further processed by a readout layer. The model parameters $\mathcal{H}$ are optimized to minimize a task-specific loss (e.g., cross-entropy for classification or mean squared error for regression tasks) over a training set. Inheriting the stability to perturbations of neural network structures (Gama et al., 2020a; Bruna & Mallat, 2013), VNNs have been shown stable to errors in covariance estimation, as stated in the following theorem.
Theorem 1 (Stability of VNNs (Sihag et al., 2022)).
Consider a VNN $\mathsf{\Phi}(\mathbf{x},\mathbf{C},\mathcal{H})$ of $L$ layers with $F$ covariance filters $\mathbf{H}(\mathbf{C})$ per layer and nonlinearities $\sigma(\cdot)$ such that $|\sigma(a)-\sigma(b)|\leq|a-b|$. Let each filter’s frequency response be Lipschitz with constant $P$ [^1] For a definition of Lipschitz filter see Def. 1.. Consider a generic data sample $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{C})$ such that $\|\mathbf{x}\|\leq 1$. Then, with probability $1-\mathcal{O}(1/t)$ it holds that

### v2

@@ Abstract
Covariance Neural Networks (VNNs) perform graph convolutions on the covariance matrix of input data to leverage correlation information as pairwise connections. They have achieved success in a multitude of applications such as neuroscience, financial forecasting, and sensor networks. However, the empirical covariance matrix on which VNNs operate typically contains spurious correlations, creating a mismatch with the actual covariance matrix that degrades VNNs’ performance and computational efficiency. To tackle this issue, we put forth Sparse coVariance Neural Networks (S-VNNs), a framework that applies sparsification techniques on the sample covariance matrix and incorporates the latter into the VNN architecture. We investigate the S-VNN when the underlying data covariance matrix is both sparse and dense. When the true covariance matrix is sparse, we propose hard and soft thresholding to improve the covariance estimation and reduce the computational cost. Instead, when the true covariance is dense, we propose a stochastic sparsification where data correlations are dropped in probability according to principled strategies. Besides performance and computation improvements, we show that S-VNNs are more stable to finite-sample covariance estimations than nominal VNNs and the analogous sparse principal component analysis. By analyzing the impact of sparsification on their behavior, we tie the S-VNN stability to the data distribution and sparsification approach. We support our theoretical findings with experimental results on a variety of application scenarios, ranging from brain data to human action recognition, and show an improved task performance, improved stability, and reduced computational time compared to alternatives.
I Introduction
@@
Covariance-based data processing is key to signal processing and machine learning pipelines due to its ability to whiten data distributions, identify principal directions, and estimate the interdependencies among features. Such advantages have been shown in several applications including estimating brain connectivity [1, 2], financial data [3, 4] and human action recognition [5, 6]. One prominent method built upon the covariance matrix is the Principal Component Analysis (PCA), which maximizes the variance of data points by projecting them onto the eigenvectors of the said matrix [7]. However, PCA-based approaches are unstable to errors in the covariance estimation, i.e., a poor estimate of the covariance matrix and its eigenvectors may lead to unpredictably bad results, especially in high-dimensional and low-data regimes or when the covariance matrix eigenvalues are close to each other [8, 9, 10]. To overcome this issue, coVariance Neural Networks (VNNs) have been proposed [11]. In essence, VNNs treat the covariance matrix as a graph with each variable as a node and the covariance value as an edge weight, and apply graph-based convolutional learning on this graph. Such an operation performs hierarchical spectral learning on the covariance graph, which ties to PCA as it learns the importance of the principal directions of the data for the task at hand. Due to the graph convolution operation, VNNs inherit the stability property of graph neural networks (GNNs) [12] and allows them to be stable to finite-data estimation errors in the covariance matrix; i.e., the output difference of a VNN operating on the sample covariance (i.e., estimated from data samples) and the true covariance is bounded [11, 13]. Additionally, they enjoy transferability across covariances of different dimensions [14]. These properties make VNNs effective in a variety of settings ranging from datasets with different resolutions [14] to temporal [15] and biased data [16]. Furthermore, VNNs represent a general framework that encompasses several methods commonly applied in practice, since estimating a graph from data correlations and applying a GNN on it is extremely common in fields such as brain connectivity estimation [1, 2], financial data processing [3, 4] and human action recognition [5, 6].
Nevertheless, VNNs operate on the sample covariance matrix, which results in two limitations. First, VNNs are sensitive to settings in which the true covariance is sparse, especially in high-dimensional and low-data regimes [9, 10, 17]. Second, VNNs are computationally expensive and memory inefficient, as the sample covariance matrix is typically dense due to estimation errors, regardless of the true covariance matrix being dense or sparse, which is an emphasized computational burden in high-dimensional datasets.
To overcome these limitations, we study the effect of sparse covariance matrix estimators on the learned representation of the VNN, and how properties of the latter affect and are affected by the sparsification strategy as well as the data distribution. More specifically, we propose Sparse coVariance Neural Networks (S-VNNs), which incorporate sparsification-based covariance regularizers into the VNN architecture, and analyze the impact of sparsification on performance when both the true covariance matrix is sparse and dense.
@@
When true covariance matrix is sparse, we incorporate thresholded covariance estimates into S-VNNs and show that they are robust to the finite-data estimation error. For the more generic scenario where the true covariance matrix is not necessarily sparse, we put forth a stochastic sparsification approach and discuss its impact on the S-VNN. Our analysis reveals an inherent trade-off between stability, sparsification degree and strategy, which, in turn, translates into a trade-off between finite-data performance and computational efficiency. We show that dropping covariances according to their absolute values leads to a higher stability but low sparsification, whereas dropping a fixed percentage of covariance values improves sparsification at the expense of stability. We corroborate our findings with experiments on synthetic and real datasets, highlighting the benefits of sparsity on covariance-based neural processing architectures in terms of both performance and computation. Our specific contributions are as follows:
(C1) Sparse covariance neural networks. We develop sparse covariance neural networks (S-VNNs) that perform sparsification techniques on the sample covariance matrix for performance improvement and computation reduction. We study the impact of sparsification on the learned embeddings and reveal a trade-off between sparsity, performance, and computational complexity.
@@
(C2) Sparsification techniques. We consider different sparsification techniques for S-VNNs. When the true covariance is sparse, we apply hard and soft thresholding; when the true covariance is dense, we propose a stochastic sparsification strategy that allows controlling the desired sparsification level in line with the found trade-off in (C1).
@@
(C3) Stability analysis. We investigate the role of different sparsification techniques and characterize the effects of covariance sparsification on the VNN performance through a finite-data sample stability analysis. The results show that S-VNNs reduce computational cost and also improve stability of the learned embeddings to covariance estimation errors, compared to the nominal VNN and the sparse PCA.
@@
(C4) Empirical validation. We validate our theoretical findings with experiments on both synthetic and real data. The results demonstrate the better performance of covariance sparsification in low-data settings and show the improved efficiency on brain data and human action recognition use cases. We make our code available [^1] https://github.com/andrea-cavallo-98/SVNN.
The rest of this paper is organized as follows. Section II introduces VNNs, their link to PCA and states our problem setting. Sections III and IV present our sparse covariance neural networks, propose different sparsification strategies when the true covariance is sparse or generic, and characterize their corresponding effects on VNN’s stability. Finally, Section V presents the experimental results.
II Preliminaries
Consider $t$ samples $\{\mathbf{x}_{i}\}_{i=1}^{t}$ of a random vector $\mathbf{x}\in\mathbb{R}^{N}$ with mean $\boldsymbol{\mu}=\mathbb{E}[\mathbf{x}]\in\mathbb{R}^{N}$ and covariance $\mathbf{C}=\mathbb{E}[(\mathbf{x}-\boldsymbol{\mu})(\mathbf{x}-\boldsymbol{\mu})^{\mathsf{T}}]\in\mathbb{R}^{N\times N}$. We can estimate these quantities from $t$ observed samples as $\boldsymbol{\hat{\mu}}=\sum_{i=1}^{t}\mathbf{x}_{i}/t$, $\mathbf{\hat{C}}=\sum_{i=1}^{t}(\mathbf{x}_{i}-\boldsymbol{\hat{\mu}})(\mathbf{x}_{i}-\boldsymbol{\hat{\mu}})^{\mathsf{T}}/t$. The estimated covariance $\mathbf{\hat{C}}$ admits the eigendecomposition $\mathbf{\hat{C}}=\mathbf{\hat{V}}\mathbf{\hat{\Lambda}}\mathbf{\hat{V}}^{\mathsf{T}}$ with orthonormal eigenvectors $\mathbf{\hat{V}}=[\mathbf{\hat{v}}_{1},\dots,\mathbf{\hat{v}}_{N}]\in\mathbb{R}^{N\times N}$ and eigenvalues $\mathbf{\hat{\Lambda}}=\text{diag}(\hat{\lambda}_{1},\dots,\hat{\lambda}_{N})$. Using the sample covariance $\mathbf{\hat{C}}$ and its eigendecomposition for data processing introduces errors related to the statistical uncertainty of the estimation, i.e., the sample covariance $\mathbf{\hat{C}}$ can significantly differ from the true covariance $\mathbf{C}$ if the number of samples $t$ is small, the dimensionality $N$ is large or the eigenvalues of the true covariance are close to each other [8]. This creates problems for techniques that model data interdependencies through their covariance, such as PCA, as we elaborate next.
II-A PCA Transform
Given a zero-mean dataset $\mathbf{X}\in\mathbb{R}^{N\times t}$ (w.l.o.g.) with each column a data sample $\mathbf{x}\in\mathbb{R}^{N}$, PCA applies a transformation $\mathbf{\tilde{X}}=\mathbf{T}^{\mathsf{T}}\mathbf{X}$ with $\mathbf{T}\in\mathbb{R}^{N\times N}$ such that the covariance of the transformed data $\mathbf{\tilde{X}}\mathbf{\tilde{X}}^{\mathsf{T}}$ is diagonal. This is achieved by setting $\mathbf{T}=\mathbf{\hat{V}}$ where $\mathbf{\hat{V}}$ is the eigenvector matrix of the covariance of the data, because $\mathbf{\tilde{X}}\mathbf{\tilde{X}}^{\mathsf{T}}=\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{X}\mathbf{X}^{\mathsf{T}}\mathbf{\hat{V}}=\mathbf{\hat{V}}^{\mathsf{T}}(\mathbf{\hat{V}}\mathbf{\hat{\Lambda}}\mathbf{\hat{V}}^{\mathsf{T}})\mathbf{\hat{V}}=\mathbf{\hat{\Lambda}}$ due to the orthonormality of the covariance eigenvectors. Since the principal directions maximize the variance of the transformed data, PCA is also used for dimensionality reduction by selecting only the $k$ covariance eigenvectors corresponding to the largest eigenvalues, i.e., $\mathbf{\tilde{X}}_{(k)}=[\mathbf{\hat{V}}]^{T}_{1,\dots,k}\mathbf{X}$ where $\cdot]_{1,\dots,k}$ selects the first $k$ columns. This operation can be interpreted as the application of an ideal spectral filter on the covariance eigenvalues $h_{\textnormal{PCA}}(\lambda)=\mathbf{1}[\lambda\geq\lambda_{k}]$, i.e., $\mathbf{\tilde{X}}_{(k)}=[\operatorname{diag}(h_{\textnormal{PCA}}(\lambda_{1}),\dots,h_{\textnormal{PCA}}(\lambda_{N}))\mathbf{\hat{V}}^{T}\mathbf{X}]_{1,\dots,k}$ – see Fig. 1.
However, PCA presents some limitations. First, operating on the eigenvectors of the sample covariance matrix $\mathbf{\hat{V}}$ makes it unstable to estimation errors, which becomes problematic for close eigenvalues and in sparse low-data regimes (e.g., when the data dimensionality is of the same order as the number of data points [18]). Second, PCA requires the explicit computation of the covariance eigendecomposition which is of order $\mathcal{O}(N^{3})$. Third, PCA is neither inductive nor transferable, i.e., small data drifts, new data samples or samples of different dimensionality require the recomputation of the principal components. Finally, PCA for dimensionality reduction focuses on the directions of largest variance, which might lose relevant information for the downstream task (e.g., if the data contains high-variance noise). To counteract these limitations, covariance filters and covariance neural networks have been proposed.
Figure 1: Frequency responses for PCA with dimensionality reduction (left) and covariance filters (right). The PCA filter selects the largest $k$ covariance eigenvalues, whereas the covariance filter learns a polynomial filter function.
II-B Covariance Filters
Covariance filters [11] are an extension of PCA that is stable to finite-sample estimation errors of the covariance matrix, even with close eigenvalues. Specifically, each of the $N$ data features is seen as a node of a graph and the data sample $\mathbf{x}\in\mathbb{R}^{N}$ as a graph signal, where the covariance $\mathbf{C}$ represents the graph structure – see Fig. 2. Covariance filters are graph convolutional filters acting on the covariance graph, i.e., they transform an input signal $\mathbf{x}$ into an output $\mathbf{u}$ as
$\mathbf{u}=\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}=\sum_{k=0}^{K}h_{k}\mathbf{\hat{C}}^{k}\mathbf{x}.$ (1)
This operation can be computed recursively, i.e., $\mathbf{C}^{k}\mathbf{x}=\mathbf{C}(\mathbf{C}^{k-1}\mathbf{x})$, with a total cost of $\mathcal{O}(kN^{2})$ which makes covariance filters more efficient than PCA. Moreover, the term $N^{2}$ is due to the density of the sample covariance. In this paper, we study sparse covariance estimators which can reduce this cost to $\mathcal{O}(k\|\mathbf{\hat{C}}\|_{0})$, where $\|\mathbf{\hat{C}}\|_{0}$ is the number of non-zero elements of the estimated covariance.
The behavior of the covariance filter can be characterized by computing the graph Fourier transform of its input and output signal, i.e., $\mathbf{\tilde{u}}=\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{u}=\mathbf{\hat{V}}^{\mathsf{T}}\sum_{k=0}^{K}h_{k}[\mathbf{\hat{V}}\mathbf{\hat{\Lambda}}\mathbf{\hat{V}}^{\mathsf{T}}]^{k}\mathbf{x}=\sum_{k=0}^{K}h_{k}\mathbf{\hat{\Lambda}}^{k}\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{x}$. For the $i$-th entry, this leads to $\mathbf{\tilde{u}}_{i}=\sum_{k=0}^{K}h_{k}\hat{\lambda}_{i}^{k}\mathbf{\tilde{x}}_{i}=h(\hat{\lambda}_{i})\mathbf{\tilde{x}}_{i}$. That is, the frequency response of the covariance filter is a polynomial $h(\lambda)=\sum_{k=0}^{K}h_{k}\lambda^{k}$ with frequency variable $\lambda$ specified on the eigenvalues $\hat{\lambda}_{i}$ – see Fig. 1. The frequency response of the covariance filter highlights its analogy with PCA, since both operations process the eigenvectors of the covariance matrix and there exists a set of coefficients $h_{k}$ such that the covariance filter recovers the operation of PCA [11, Thm. 1]. Based on the frequency response, we introduce the definition of Lipschitz covariance filter, which will be instrumental in the following analysis.
Definition 1 (Lipschitz covariance filter).
The covariance filter is Lipschitz with constant $P$ if, for every pair of eigenvalues $\lambda_{i},\lambda_{j}\in[0,\lambda_{\textnormal{max}}]$, $\lambda_{i}\neq\lambda_{j}$, its frequency response satisfies: $|h(\lambda_{i})-h(\lambda_{j})|\leq P|\lambda_{i}-\lambda_{j}|$, where $\lambda_{\textnormal{max}}\in[0,\infty)$ identifies a suitable range for the covariance eigenvalues.
Lipschitz covariance filters control the variability of the frequency response through the constant $P$. Inheriting the stability of graph convolutional filters [12], covariance filters have been shown stable to errors in covariance estimation. To report this result, we introduce the following definitions:
$k_{j}=(\mathbb{E}[\|\mathbf{x}\mathbf{x}^{\mathsf{T}}\mathbf{v}_{j}\|^{2}]-\lambda_{j}^{2})^{1/2},\quad k_{\textnormal{max}}=\max_{j}k_{j},$ (2)
$k_{\textnormal{min}}=\min_{j,\lambda_{j}>0}k_{j},\quad\kappa=\max_{i,j:\lambda_{i}\neq\lambda_{j}}k_{i}^{2}/|\lambda_{i}-\lambda_{j}|,$ (3)
which are terms that relate to the kurtosis of the data distribution (we refer to [11] for additional details as these are not part of our subsequent analysis). The stability of covariance filters is stated next.
Theorem 1 ([11]).
Consider a covariance filter $\mathbf{H}(\mathbf{C})$ [cf. (1)] that is Lipschitz with constant $P$ [Def.˜1]. Consider a true covariance $\mathbf{C}$ and its sample estimate from $t$ samples $\mathbf{\hat{C}}$, respectively. With probability at least $1-t^{-2\epsilon}-2\kappa N/t$, for any $\epsilon\in(0,1/2]$ it holds that